A bicycle trip of 120 miles would have taken 3 hours less if the average speed had been increased by 2 miles per hour. Find the average speed of the bicycle.
Found 2 solutions by checkley77, themathprof:
Answer by checkley77(12844)

120/X=[120/(X+2)]+3
X[120/(X+2)+3=120
120X/(X+2)+3X=120
[(120X+3X(X+2)]/(X+2)=120
(120X+3X^2+6X)/(X+2)=120
126X+3X^2=120(X+2)
3X^2+126X=120X+240
3X^2+126X-120X-240=0
3X^2+6X-240=0
3(X^2+2X-80)=0
3(X+10)(X-8)=0
X-8=0
X=8 MPH IS THE SLOWER TRIP.
8+2=10 MPH FOR THE FASTER TRIP.
120/8=15 HOURS.
120/10=12 HOURS.
240/X=27
27X=240
X=240/27
X=8.89 MPH IS THE AVERAGE SPEED.



Answer by themathprof(21)

A bicycle trip of 120 miles would have taken 3 hours less if the average speed had been increased by 2 miles per hour. Find the average speed of the bicycle.
x= avg speed of bicycle
The distance is the same (120) before and after the increase
The Rate changes from x to x+2
Note that in each case the time is Distance/Rate

Rate x Time = Distance
Before increase
x....120/x.....120
After increase
x+2...120/x+2..120

New Time is 3 less
120/(x+2) = 120/x -3
Multiply each term by LCM=x(x+2)
x(x+2)(120)/(x+2) = x(x+2)(120)/x -3(x)(x+2)
Cancel denominators
120x=120(x+2)-3x(x+2)
120x=120x+240-3x^2-6x
0=-3x^2-6x+240
divide by -3
0=x^2+2x-80
Factor right side
0=(x-8)(x+10)
x-8=0 .... x=8 avg speed of bicycle
x+10=0 ... x=-10 (reject, speed can't be negative!)

CHECK
Rate x Time = Distance
Before increase
8....120/8=15.....120
After increase
8+2=10...120/8+2=12..120
Time decreases 3 from 15 to 12
See more word problems worked out step by step
in my math911 free tutorial software.
Download, install, activate in minutes.
www.math911.com
No Charge! Tell your classmates, friends,
(and teacher)
Prof Martin Weissman
mathprof@hotmail.com